A traffic study of 2,500 vehicles that passed by a checkpoint showed that their speeds were normally distributed with a mean of 82.5 kph and a standard deviation of 11.5 kph. How many vehicles had a speed of more than 90 kph? *


1
Expert's answer
2021-11-12T01:49:44-0500

Solution:

"\bar x = 82.5"

"s = 11.5"

"P(X>90)=1-P(X\le90)=1-P(Z\le\frac{90-\bar x}{s})="

"=1-P(Z<\frac{90-82.5}{11.5})=1-P(Z<0.65)=1-0.74215=0.25785=25.785\%"

25.785% of 2500 vehicles = "2500 \times 0.25785=644.625\approx645"

Thus, 645 vehicles have a speed of more than 90kph.